# Frostman random variables, entropy inequalities, and applications

### v3

@@ 4.2 Diagonal forms with dependent random variables
On the other hand, We can apply Lemma A.2, then combine (45) to get that
$\begin{split}&\frac{H_{n}(X_{1})+H(Y_{1})}{2}\leq\frac{H_{n}(X)+H_{n}(Y)}{2P(E)}\leq(1+O(C\eta^{2s}))(ns+\Psi),\\ &H_{n}(X_{1}+Y_{1})\leq\frac{H_{n}(X+Y)}{P(E)}\leq(1+O(C\eta^{2s}))(ns+\Psi),\\ &H_{n}(\alpha X_{1}^{2}+\beta Y_{1}^{2})\leq\frac{H_{n}(\alpha X^{2}+\beta Y^{2})}{P(E)}\leq(1+O(C\eta^{2s}))(ns+\Psi).\end{split}$ (47)
@@
$\begin{split}&H_{n}(X_{1})+H_{n}(Y_{1})-H_{n}(X_{1},Y_{1})=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1),\\ &H_{n}(X_{1}+Y_{1})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1),\\ &H_{n}(\alpha X_{1}^{2}+\beta Y_{1}^{2})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1).\end{split}$
Now we write $\log K:=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1)$. Applying Theorem 3.1 with the random variable $(X_{1},Y_{1})$, we obtain the random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$, where $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$, and
$H_{n}(X^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(X_{1})-\log K,\quad H_{n}(Y^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(Y_{1})-\log K.\\$ (48)
First, let us take the functions $f$ and $g$ in Theorem 3.1 to be the identity functions. Then
$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}\leq 7\log K+O(1).$ (49)
Second, we take the functions $f,g$ in Theorem 3.1 to be $f(x)=\alpha x^{2}$ and $g(y)=\beta y^{2}$. By the construction $(X_{1},Y_{1})\equiv(X,Y||(X,Y)|>\eta)$, the ranges of $X^{\prime\prime}$ and $Y^{\prime\prime}$ are contained in $c,-\eta/\sqrt{2}]\cup[\eta/\sqrt{2},c]$ with probability $1$. Now, by Example 2.12, the random variables $f(X^{\prime\prime})$ and $g(Y^{\prime\prime})$ are weakly $O_{\phi}(\max\{\eta^{-1},c\})$-bi-Lipschitz functions of $X^{\prime\prime}$ and $Y^{\prime\prime}$, respectively. It follows that
$H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|X^{\prime},Y^{\prime})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}\leq 7\log K+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi}(1).$ (50)
@@
$\qquad=\frac{P_{E}(X_{1}\in B(z,r),D_{n}(Y_{1})=y)}{P_{E}(D_{n}(Y^{\prime})=y)}\leq\frac{2P(X\in B(z,r),D_{n}(Y)=y)}{P_{E}(D_{n}(Y^{\prime})=y)}.$
@@
This completes the proof for the claims given in (51). We mention that these two Frostman constants in (51) are $\gg 1$.
@@
As the expressions to follow get rather lengthy, we now introduce some shorthand. In what follows, $H_{n}(~\cdot~|x,y)$ is an abbreviation of $H(D_{n}(\cdot)~|D_{n}(X^{\prime},Y^{\prime})=(x,y)).$ By Lemma A.2, there exist random variables, depending on $(x,y)$, that have the same distributions with $(X^{\prime\prime},Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y))$, respectively. These random variables take the same value as $(X^{\prime\prime},Y^{\prime\prime})$ does. By construction, their ranges are contained in $\big{(}[-c,-\eta/\sqrt{2}]\cup[\eta/\sqrt{2},c]\big{)}^{2}$ with probability $1$. Also note that $(X^{\prime\prime}|D_{n}(X_{1},Y_{1})=(x,y))$ and $(Y^{\prime\prime}|D_{n}(X_{1},Y_{1})=(x,y))$ are independent. Recall that the constant $\epsilon_{1}=\epsilon_{1}(s)$ in Lemma 4.1 is independent of the concrete random variables. Finally, recall (51). Now we can apply Lemma 4.1 with these random variables, with $C_{1}=O\big{(}C\cdot 2^{-ns_{2}}/P_{E}(D_{n}(Y^{\prime})=y)\big{)}$, $C_{2}=O\big{(}C\cdot 2^{-ns_{1}}/P_{E}(D_{n}(X^{\prime})=x)\big{)}$, $c_{1}=c$ and $c_{2}=\eta$, to deduce that
$\max\left\{H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y),H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)\right\}\geq n(s+\epsilon_{1})$
$\qquad+O\big{(}\log_{+}C-n(s_{1}+s_{2})-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big{)}$
$\qquad+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1)$ (52)
for some $\epsilon_{1}=\epsilon_{1}(s)>0$. By conditional independence, and combining (13), we also have that
$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)\geq\max\{H_{n}(X^{\prime\prime}|x,y),H_{n}(Y^{\prime\prime}|x,y)\}+O(1).$
Using a similar argument, Lemma A.2, and Example 2.16, one obtains
$H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)\geq\max\{H_{n}(\alpha X^{\prime\prime 2}|x,y),H_{n}(\beta Y^{\prime\prime 2}|x,y)\}+O(1)$
$\qquad=\max\{H_{n}(X^{\prime\prime}|x,y),H_{n}(Y^{\prime\prime}|x,y)\}+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi}(1).$
Then
$\min\{H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y),H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)\}$
$\qquad\geq\frac{H_{n}(X^{\prime\prime}|x,y)+H_{n}(Y^{\prime\prime}|x,y)}{2}+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi}(1).$ (53)
Therefore given the condition that $D_{n}(X_{1},Y_{1})=(x,y)$, it follows from (52) and (53) that
$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)+H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)-\frac{H_{n}(X^{\prime\prime}|x,y)+H_{n}(Y^{\prime\prime}|x,y)}{2}$
$\quad\geq n(s+\epsilon_{1})+O\big{(}-n(s_{1}+s_{2})-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big{)}$
$\qquad\qquad\qquad\quad+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$
@@
Now we sum all of these expressions respective to each $(x,y)$ with the weight $P_{E}(D_{n}(X^{\prime},Y^{\prime})=(x,y))$. Note that
$\sum\limits_{x,y}P_{E}(D_{n}(X^{\prime},Y^{\prime})=(x,y))H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)$
$\qquad=\sum\limits_{x,y}P_{E}(D_{n}(X^{\prime\prime},Y^{\prime\prime})=(x,y))H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)=H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime}),\text{ and}$
$-\sum\limits_{x,y}P_{E}(D_{n}(X^{\prime},Y^{\prime})=(x,y))\log P_{E}(D_{n}(X^{\prime})=x)$
$\qquad=-\sum\limits_{x}P_{E}(D_{n}(X^{\prime})=x)\log P_{E}(D_{n}(X^{\prime})=x)=H_{n}(X^{\prime}).$
The other terms behave similarly. Therefore,
$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|X^{\prime},Y^{\prime})-\frac{H_{n}(X^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(Y^{\prime\prime}|X^{\prime},Y^{\prime})}{2}$
$\qquad\qquad\geq n(s+\epsilon_{1})+O(-n(s_{1}+s_{2})+H_{n}(X^{\prime})+H_{n}(Y^{\prime}))$
$\qquad\qquad\qquad\quad+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$ (54)
@@
Substituting from (48)–(50), and further combining with $(X^{\prime},Y^{\prime})\equiv(X_{1},Y_{1})$ we obtain that
$\begin{split}&\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}+7\log K\right)+\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}+7\log K\right)\\ &\qquad-\frac{(H_{n}(X_{1})-\log K)+(H_{n}(Y_{1})-\log K)}{2}-ns+O(H_{n}(X_{1})+H_{n}(Y_{1})-n(s_{1}+s_{2}))\\ &\qquad\geq n\epsilon_{1}+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).\end{split}$ (55)
@@
Recalling (47), the fact $s_{1}+s_{2}\geq 2s$, and the notation $\log K=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1)$, the left-hand side of (55) is
$\leq O\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}-ns\right)+15\log K$
$=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$
@@
So there is some absolute constant $\tilde{C}>0$ such that
$\tilde{C}\Psi\geq n\epsilon_{1}+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$
Recall that we require $0<\eta<\min\{2^{-1}(1/2C)^{1/2s},1/10\}$, $0<\epsilon_{1}<1$, and $\epsilon_{1}$ only depends on $s$. Now assume that the implied constant in the term $O(C\eta^{2s}n)$ is $C^{\prime}$, with $C^{\prime}\geq 1$. We then take $\eta=\min\Big{\{}\frac{1}{2}\big{(}\frac{\epsilon_{1}}{2C^{\prime}C}\big{)}^{1/2s},\,\frac{1}{10}\Big{\}}$. Then $C^{\prime}C\eta^{2s}n\leq n\epsilon_{1}/2$. Therefore,
$\Psi\gg n\epsilon_{1}+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{\phi,s}(1).$
The proof is completed. ∎
4.3 The general forms
Any non-degnerate rational quadratic from $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ can be diagonalized by rational linear change of variables, i.e.,
$\phi(x,y)=\alpha(\xi_{11}x+\xi_{12}y)^{2}+\beta(\xi_{21}x+\xi_{22}y)^{2}$ (56)
for some $\alpha,\beta\in\mathbb{Q}^{\ast}$ and $\xi_{11},\xi_{12},\xi_{21},\xi_{22}\in\mathbb{Q}$ with $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$.
One can verify that $\phi(x,y)$ is diagonal if and only if exactly two of the coefficients $\xi_{11},\xi_{12},\xi_{21},\xi_{22}$ are zero. For non-diagonal case, at most one of $\xi_{11},\xi_{12},\xi_{21},\xi_{22}$ may be zero. In the following, we will show that, if exactly one of these numbers is zero, then we can transform $\phi(x,y)$ into another form like (56) with all corresponding coefficients non-zero. Now we assume without loss of generality that $\xi_{11},\xi_{21},\xi_{22}\neq 0$ and $\xi_{21}=0$. By dividing $\xi_{11}$ and $\xi_{12}$ in the two brackets, respectively, we may rewrite $\phi(x,y)=\alpha^{\prime}x^{2}+\beta^{\prime}(x+d^{\prime}y)^{2}$, with $\alpha^{\prime},\beta^{\prime},d^{\prime}\in\mathbb{Q}^{\ast}$. Now, let us take some $a,c\in\mathbb{Q}^{\ast}$ satisfying $a\neq c$, $a\neq 1/d^{\prime}$, $c\neq 1/d^{\prime}$ and
$(d^{\prime}a-1)(d^{\prime}c-1)=-\alpha^{\prime}/\beta^{\prime}.$
One can check that
$\phi(x,y)=\frac{\beta^{\prime}d^{\prime}(1-cd^{\prime})}{a-c}(ax+y)^{2}+\frac{\beta^{\prime}d^{\prime}(ad^{\prime}-1)}{a-c}(cx+y)^{2}.$
All coefficient in above formula are non-zero rational numbers. Furthermore, we can assume in (56) without loss of generality that $\xi_{11}\neq-\xi_{21}$ and $\xi_{12}\neq-\xi_{22}$, because non-zero constants can be multiplied inside both brackets. In all situations, all the coefficients depend on $a_{1},a_{2},a_{3}$.
@@ 4.3 The general forms
Theorem 1.3’. Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C_{1},C_{2},c>0$. There is a positive constant $\epsilon_{3}=\epsilon_{3}(s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon_{3})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{s}(1).$
@@
Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.30 it is also jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman. Applying Theorem 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(s)$ such that
$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}$
$\qquad\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{s}(1)$
for all $n\in\mathbb{N}$. In view of $s_{1}+s_{2}\geq 2s$, Lemma 2.32(ii) shows that
$H_{n}(X,Y)\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+ns+O(\log_{+}C_{1}C_{2})+O(1).$
@@
$\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$.
Theorem 1.3”. Suppose that $\phi(x,y)$ has the form (56), where $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $\xi_{11},\xi_{12},\xi_{13},\xi_{14}\in\mathbb{Q}^{\ast}$, $\xi_{11}\neq-\xi_{21}$, $\xi_{12}\neq-\xi_{22}$ and $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$. Let $0<s<1$ and $C_{1},C_{2},c>0$. There is a positive constants $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1).$
@@
Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, Lemma 2.30 tells that it is both marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman and jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman. It follows from Lemma 2.32(i) and the fact $s_{1}+s_{2}\geq 2s$ that
$\min\big{\{}H_{n}(X)+H_{n}(Y),\,H_{n}(X,Y)\big{\}}\geq 2ns-\log C_{1}C_{2}+O(1).$ (57)
@@
$H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\leq H_{n}(X,Y)-2ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).$ (58)
Consider the random variable $(\xi_{11}X+\xi_{12}Y,\,\xi_{21}X+\xi_{22}Y)$. Its range is contained in $c^{\prime\prime},c^{\prime\prime}]^{2}$, with $c^{\prime\prime}=O_{\phi}(c)$. And Lemma 2.31 shows that it is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, with $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$, since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. By Lemma 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(s)>0$ such that
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$ (59)
@@
Notice that $\xi_{11}\neq-\xi_{21}$ and $\xi_{12}\neq-\xi_{22}$. It follows from Theorem 2.23 that
@@
for some $d_{1}^{\prime}=d_{1}^{\prime}(\phi)>0$ and $d_{2}^{\prime}=d_{2}^{\prime}(\phi)\geq 0$. And also
@@
for some $d_{1}^{\prime\prime}=d_{1}^{\prime\prime}(\phi)>0$ and $d_{2}^{\prime\prime}=d_{2}^{\prime\prime}(\phi)\geq 0$.
@@
In the following, Theorem 1.4 will be proved in the following form.
Theorem 1.4’ Let $\phi(x,y)$ be in the form (56). Let $0<s<1$, $C_{1},C_{2},c>0$. There is a positive constant $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that the following statement holds: Let $X$ and $Y$ be independent real random variables with ranges contained in [c,c][-c,c]. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$

### v4

@@ 4.3 Diagonal polynomials with dependent random variables
On the other hand, We can apply Lemma A.2 with $P(E)^{-1}\leq 1+\epsilon_{1}^{\prime\prime}/\mathfrak{r}$, then combine (64) to get that
$\frac{H_{n}(X_{0})+H(Y_{0})}{2},\,H_{n}(X_{0}+Y_{0}),\,H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))\,\leq(1+\epsilon_{1}^{\prime\prime}/\mathfrak{r})(ns+\Psi).$ (67)
@@
$\begin{split}&H_{n}(X_{0})+H_{n}(Y_{0})-H_{n}(X_{0},Y_{0})\leq\log K\\ &H_{n}(X_{0}+Y_{0})-\frac{H_{n}(X_{0})+H_{n}(Y_{0})}{2}\leq\log K,\\ &H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))-\frac{H_{n}(X_{0})+H_{n}(Y_{0})}{2}\leq\log K,\end{split}$
where $\log K=O(\Psi)+O(n\epsilon^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K})$.
We now apply Theorem 3.1 to $(X_{0},Y_{0})$, and obtain a quadruple $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ with certain properties (see (i)-(iv) in Theorem 3.1). In particular, the random variables $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$. Let us take the functions $f$ and $g$ in Theorem 3.1 to be the identity functions. Then
$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})-\frac{H_{n}(X_{0})+H_{n}(Y_{0})}{2}\leq 7\log K+O(1).$ (68)
Again, we take the functions $f,g$ in Theorem 3.1 to be $f(x)=\rho_{1}(x)$ and $g(y)=\rho_{2}(y)$. With the weakly $(\kappa,d)$-bi-Lipschitz property, it satisfies that
$H_{n}(\rho_{1}(X^{\prime\prime})+\rho_{2}(Y^{\prime\prime})|X^{\prime},Y^{\prime})-\frac{H_{n}(X_{0})+H_{n}(Y_{0})}{2}\leq 7\log K+O(\log\mathfrak{K}).$ (69)
@@
$\qquad=\frac{P_{E}(X_{0}\in B(z,r),D_{n}(Y_{0})=y)}{P_{E}(D_{n}(Y^{\prime})=y)}\leq\frac{2P(X\in B(z,r),D_{n}(Y)=y)}{P_{E}(D_{n}(Y^{\prime})=y)}.$
@@
This completes the proof for the claims given in (70).
@@
As the expressions to follow get rather lengthy, we now introduce some shorthand. In what follows, $H_{n}(~\cdot~|x,y)$ is an abbreviation of $H(D_{n}(\cdot)~|D_{n}(X^{\prime},Y^{\prime})=(x,y)).$ By Lemma A.2, there exist random variables, depending on $(x,y)$, that have the same distributions with $(X^{\prime\prime},Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y))$, respectively. Similar as previously, we can apply Lemma 4.8 with these random variables, with $C_{1}=O\big(C\cdot 2^{-ns_{2}}/P_{E}(D_{n}(Y^{\prime})=y)\big)$ and $C_{2}=O\big(C\cdot 2^{-ns_{1}}/P_{E}(D_{n}(X^{\prime})=x)\big)$, to deduce that
$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)+H_{n}(\rho_{1}(X^{\prime\prime})+\rho_{2}(Y^{\prime\prime})|x,y)\geq n(2s+\epsilon_{1}^{\prime\prime})+O_{d}(\log_{+}c)+O_{\phi,s}(1)$
$\qquad+O_{d,s}\big(\log_{+}C-n(s_{1}+s_{2})-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big).$
@@
Now we sum all of these expressions respective to each $(x,y)$ with the weight $P_{E}(D_{n}(X^{\prime},Y^{\prime})=(x,y))$, which yields
$\begin{split}&H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(\rho_{1}(X^{\prime\prime})+\rho_{2}(Y^{\prime\prime 2})|X^{\prime},Y^{\prime})\geq n(2s+\epsilon_{1}^{\prime\prime})\\ &\qquad\qquad\qquad+O_{d,s}(-n(s_{1}+s_{2})+H_{n}(X^{\prime})+H_{n}(Y^{\prime}))+O(\log\mathfrak{K}).\end{split}$ (71)
@@
Substituting from (66), (68),(69), and further combining with $(X^{\prime},Y^{\prime})\equiv(X_{0},Y_{0})$ we obtain that
$H_{n}(X_{0})+H_{n}(Y_{0})+14\log K+O(\log\mathfrak{K})\geq n(2s+\epsilon_{1}^{\prime\prime})$ (72)
@@
Recalling (67) and the notation $\log K=O(\Psi)+O(n\epsilon^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K})$, the left-hand side of (72) is
$\leq 2ns+O(\Psi)+O(n\epsilon^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K}).$ (73)
@@
Combining (72) and (73), there is some absolute constant $\tilde{C}>0$ such that
$\tilde{C}\Psi\geq n\epsilon_{1}^{\prime\prime}+O(\log\mathfrak{K})+O(n\epsilon^{\prime\prime}/\mathfrak{r}).$
We now take $\mathfrak{r}\geq 1$ to be some absolute constant such that the last term on the right-hand side of above formula satisfies $O(n\epsilon_{1}^{\prime\prime}/\mathfrak{r})\leq n\epsilon_{1}^{\prime\prime}/2$, Therefore,
$\Psi\geq n\epsilon_{2}+O(\log\mathfrak{K})$
for $\epsilon_{2}=\epsilon_{1}^{\prime\prime}/2\tilde{C}$. The proof is completed. ∎
4.4 More general polynomials
@@
Theorem 1.3’. Let $d\geq 2$ and $\rho_{1}$, $\rho_{2}$ be real polynomials of degree $d$. Let $0<s<1$ and $C_{1},C_{2},c>0$. There is a positive constant $\epsilon_{3}=\epsilon_{3}(d,s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{3})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\rho_{1},\rho_{2},s}(1).\end{split}$
@@
Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.30 it is also jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman and marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. Applying Theorem 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(d,s)$ such that
$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\right\}$
$\qquad\qquad\geq n(s+\epsilon_{2})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\rho_{1},\rho_{2},s}(1)$
for all $n\in\mathbb{N}$. Here the implied constant in the term $O(\log_{+}c)$ is independent of $s$, since the marginally Frostman constants do not involve $\max\{1,c\}$ as in (65). In view of $s_{1}+s_{2}\geq 2s$, Lemma 2.32(ii) shows that
$H_{n}(X,Y)\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+ns-\frac{1}{2}\log C_{1}C_{2}+O(1).$
@@
Theorem 1.3”. Suppose that $\phi(x,y)$ has the form (1). Let $0<s<1$ and $C_{1},C_{2},c>0$. There is a positive constants $\epsilon_{4}=\epsilon_{4}(d,\Xi,s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{4})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1).\end{split}$
@@
Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, Lemma 2.30 tells that it is both marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman and jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman. Then
$H_{n}(X)+H_{n}(Y)\geq H_{n}(X,Y)\geq 2ns-\log C_{1}C_{2}+O(1).$ (74)
@@
$H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\leq H_{n}(X,Y)-2ns+\log C_{1}C_{2}+O(1).$ (75)
Consider the random variable $(\xi_{11}X+\xi_{12}Y,\,\xi_{21}X+\xi_{22}Y)$. Its range is contained in $c^{\prime\prime},c^{\prime\prime}]^{2}$, with $c^{\prime\prime}=O_{\phi}(c)$. And Lemma 2.31, we assume without loss of generality that it is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, with $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$, since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. By Lemma 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(d,s)>0$ such that
@@
$\qquad\qquad\qquad\qquad\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}+\log_{+}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$ (76)
@@
By Theorem 2.23, we have
@@
for some $d_{1}^{\prime}=d_{1}^{\prime}(\Xi)>0$ and $d_{2}^{\prime}=d_{2}^{\prime}(\Xi)\geq 0$. And also
@@
for some $d_{1}^{\prime\prime}=d_{1}^{\prime\prime}(\Xi)>0$ and $d_{2}^{\prime\prime}=d_{2}^{\prime\prime}(\Xi)\geq 0$.
@@
For Theorem 1.4, the situation that $\phi(x,y)$ is diagonal has been proved by Lemma 4.8. The general case will be proved in the following form.
Theorem 1.4’ Suppose that $\phi(x,y)$ has the form (1). Let $0<s<1$, $C_{1},C_{2},c>0$. There is a positive constant $\epsilon_{4}=\epsilon_{4}(d,\Xi,s)$ such that the following statement holds: Let $X$ and $Y$ be two independent real random variables with ranges contained in [c,c][-c,c]. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{d,s}(\log_{+}C_{1}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1)$